OpenStudy (anonymous):

A washer and a dryer cost $624 combined. The washer costs $76 less than the dryer. What is the cost of the dryer?

OpenStudy (anonymous):

x + x - 76 = 624 2x = 700 x = 350 dryer costs $350

OpenStudy (lalaly):

w=washer price d=dryer price w+d=624 w=d-76 so (d-76)+d=624 2d=700 d=350$

hero (hero):

w + d = 624 w = d - 76 w = 624 - d w = w d - 76 = 624 - d 2d = 700 d = 350 w = 274 So, the cost of the dryer is $350